OpenStudy (anonymous):

Evaluate 2sin(pi/3)cos(pi/3)

OpenStudy (anonymous):

@jim_thompson5910 Hii are you busy? Up for any IB SL Math questions tonight?

OpenStudy (freckles):

you could use unit circle to evaluate both sin(pi/3) and cos(pi/3)

OpenStudy (freckles):

you could also use the double angle identity and use unit circle once

OpenStudy (anonymous):

Yes, I have it in front of me but I'm confused on what to do

OpenStudy (freckles):

can you find pi/3 on the unit circle?

OpenStudy (anonymous):

yea it's 60 degrees right?

OpenStudy (freckles):

there should be an ordered pair designated to it

OpenStudy (freckles):

(x,y)=(cos(),sin())

OpenStudy (anonymous):

1/2, sqrt3/2

OpenStudy (freckles):

so that means the cos(pi/3) is 1/2 and sin(pi/3) is sqrt(3)/2 plug in

OpenStudy (anonymous):

ok gimme a sec

OpenStudy (anonymous):

ok, just for clarification, what exactly am I doing once I plug them in? I mean am I just dividing, multiplying, etc? By the way, the answer in the back is (0.766, -0.643)

OpenStudy (anonymous):

uhh hello..?

OpenStudy (freckles):

2sin(pi/3)cos(pi/3) is multiplication

OpenStudy (freckles):

abc means a times b times c

OpenStudy (anonymous):

Ohh ok my bad. so i got .037

OpenStudy (freckles):

how did you get that?

OpenStudy (anonymous):

I just plugged into my calculator 2sin(pi/3) * cos(pi/3)

OpenStudy (anonymous):

its not right, i know ;-;

OpenStudy (freckles):

sin(pi/3) is sqrt(3)/2 so replace sin(pi/3) with sqrt(3)/2 cos(pi/3) is 1/2 so replace cos(pi/3) with 1/2 \[2[\sin(\frac{\pi}{3})][\sin(\frac{\pi}{3})] \\ 2[ \frac{\sqrt{3}}{2}][\frac{1}{2}] = \frac{2}{2} [\sqrt{3}][\frac{1}{2}]=....\] you can simplify a little

OpenStudy (freckles):

oops one of those sin were suppose to read cos

OpenStudy (anonymous):

the second one right?

OpenStudy (usukidoll):

\[2[\sin(\frac{\pi}{3})][\cos(\frac{\pi}{3})] \\ 2[ \frac{\sqrt{3}}{2}][\frac{1}{2}] = \frac{2}{2} [\sqrt{3}][\frac{1}{2}]=....\]

OpenStudy (anonymous):

ahh ok so did you get rid of the 2 in the beginning and then add in the 2/2 as a separate thing?

OpenStudy (freckles):

multiplication is commutative ab =ba

OpenStudy (freckles):

\[2 \cdot \frac{\sqrt{3}}{2} \cdot \frac{1}{2} \\ =\frac{2}{1} \cdot \frac{\sqrt{3}}{2} \cdot \frac{1}{2} \\ =\frac{2 }{2 } \cdot \frac{\sqrt{3}}{1} \cdot \frac{1}{2}\]

OpenStudy (anonymous):

gotcha. But at any rate, I ended up getting 4.71

OpenStudy (anonymous):

which, again, doesn't seem right since the answer in the back of the book is totally different

OpenStudy (freckles):

you do know 2/2=?

OpenStudy (anonymous):

1

OpenStudy (freckles):

right so you have \[\sqrt{3} \cdot \frac{1}{2} \text{ which is just } \frac{\sqrt{3}}{2}\] I don't see how you are getting a number bigger than 1 ...

OpenStudy (freckles):

since sqrt(3)/2 is definitely less than 1

OpenStudy (anonymous):

i think it's just cos im plugging it into my scientific calc then

OpenStudy (freckles):

ok

OpenStudy (anonymous):

oh wait im so dumb i was doing pi3

OpenStudy (anonymous):

im so sorry my bad

OpenStudy (anonymous):

its .866 right

OpenStudy (freckles):

yes sqrt(3)/2 is approximately .866

OpenStudy (usukidoll):

yeah

OpenStudy (anonymous):

ok but the answer says it's .766, -0.643

jimthompson5910 (jim_thompson5910):

@brriiarr the fact that it shows two answers, when it should be one answer, suggests that maybe there's a mixup between the problem and answer

OpenStudy (freckles):

I don't know what your book is talking about the answer is sqrt(3)/2 which could be written approximately as .866

OpenStudy (usukidoll):

maybe it's a typo for .766 ???

OpenStudy (usukidoll):

Ah yes, sometimes there are typos in book. No book is published perfectly. sigh

OpenStudy (anonymous):

crap ok. it says in a weird little font next to it sqrt3/2 but i guess you guys were right. umm..do you think i could ask for help on one more?

OpenStudy (freckles):

you might also want to check the question you type was the question you meant and the answer that correspond to the question you asked was the right correspondence you chose

OpenStudy (anonymous):

i think it really was just a typo :p

OpenStudy (anonymous):

um but yea I have on more i needed a little help on

OpenStudy (anonymous):

ok lol so it's just to find the perimeter and area of the sector. I'll draw it

OpenStudy (anonymous):

rty

OpenStudy (anonymous):

ahh so yea, i dunno how to do this at all